OpenStudy (cutestgirlaround):

A pair of shoes usually sells for $70. If the shoes are 20% off, and sales tax is 5%, what is the total price of the shoes, including tax? $59.50 $55.12 $58.80 $62.48

OpenStudy (cutestgirlaround):

@dan815 @KendrickLamar2014

OpenStudy (cutestgirlaround):

@nikeboi101 @daisyduck04

OpenStudy (kendricklamar2014):

Lets go step by step....

OpenStudy (kendricklamar2014):

\[$70 - 20 ~percent = ?\] @cutestgirlaround

OpenStudy (cutestgirlaround):

I failed si it doesnt matter

OpenStudy (kendricklamar2014):

Oh ok, well let me show you how to answer this: \[.20 \times 70 = 14\] \[$70 - $14 = $56\] \[$56 + 5 ~ percent = $58.80 \] \[Answer: $58.80\]

OpenStudy (daisyduck04):

sorry i dont know